OpenStudy (anonymous):

The length of a rectangle is four times its width. If the perimeter is at most 130 centimeters, what is the greatest possible value for the width? Write an inequality to model the problem.

OpenStudy (haichi):

4w=130

OpenStudy (haichi):

The width is 13

OpenStudy (haichi):

13*4=52 52+52=104 13 +13 for the width is 26 so 104+26=130